OpenStudy (anonymous):

Darren designed this triangular shaped entrance, ABC, to a mall. Which expression can be used to find the distance between the points B and C of the entrance?

OpenStudy (anonymous):

A. 15 /cot 65 +15 / sec 25 B. 15 / cot 65 + 15 cot 25 C. 15/ sec 65 + 15 sec 25 D. 15 /cot 65 + 15/cot 25

jimthompson5910 (jim_thompson5910):

BC = BD + DC so you need to find BD and DC I'll help you find BD tan(angle) = opp/adj tan(65) = BD/AD tan(65) = BD/15 15*tan(65) = BD BD = 15*tan(65)

jimthompson5910 (jim_thompson5910):

Now because there isn't a tangent expression anywhere in the answer choices, but there is a cotangent expression, this means we must say BD = 15*tan(65) BD = 15*(1/cot(65) ) BD = 15/cot(65)

jimthompson5910 (jim_thompson5910):

Using the fact that tan(x) = 1/cot(x)

jimthompson5910 (jim_thompson5910):

How would you find DC?

OpenStudy (anonymous):

Do I use the reciproocal form ...

jimthompson5910 (jim_thompson5910):

Use whatever form you are familiar with and we'll go from there

OpenStudy (anonymous):

hypotenuse over adjacent or adjacent over opposite ..

jimthompson5910 (jim_thompson5910):

AD is given to be 15 ft it is opposite the given angle of 25 degrees so AD is the opposite side

jimthompson5910 (jim_thompson5910):

we want DC, which is the adjacent side

jimthompson5910 (jim_thompson5910):

so what trig function uses opposite and adjacent?

OpenStudy (anonymous):

cotangent

jimthompson5910 (jim_thompson5910):

good, cot = adj/opp

jimthompson5910 (jim_thompson5910):

so cot(25) = DC/AD cot(25) = DC/15 15*cot(25) = DC CD = 15*cot(25)

jimthompson5910 (jim_thompson5910):

DC = 15*cot(25) i meant to write

jimthompson5910 (jim_thompson5910):

now go back to the breakdown of BC BC = BD + DC and plug in what we just found BC =15/cot(65) + 15*cot(25)

OpenStudy (anonymous):

that leave us with choice B or D

OpenStudy (anonymous):

How do you know weither to just write it out like that or in fraction from.

jimthompson5910 (jim_thompson5910):

there's a slight difference

jimthompson5910 (jim_thompson5910):

I just followed the trig definitions and solved for the unknowns

jimthompson5910 (jim_thompson5910):

then plugged them back into BC = BD + DC

OpenStudy (anonymous):

Thank You. (:

OpenStudy (anonymous):

Do you mind to help me with one more questioned I have it worked out just don't know weither to write it out or put it in fraction from ?

jimthompson5910 (jim_thompson5910):

sure go ahead and post tell me what answer you went with

OpenStudy (anonymous):

The sail of a boat is in the shape of a right triangle, as shown below Which expression shows the length, in meters, of the base of the sail?

OpenStudy (anonymous):

7 cossec 50 or 7/cossec 50

jimthompson5910 (jim_thompson5910):

let x = length of base csc(50) = hyp/opp csc(50) = 7/x x*csc(50) = 7 x = 7/csc(50)

jimthompson5910 (jim_thompson5910):

That's not the same as 7*csc(50) because we're multiplying here instead of dividing

OpenStudy (anonymous):

So it is in fraction form

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thank you lots. I would give you medals but iM not sure how to give them

jimthompson5910 (jim_thompson5910):

that's ok, glad to be of help

OpenStudy (anonymous):

(: